# NuSegDG: Integration of Heterogeneous Space and Gaussian Kernel for Domain-Generalized Nuclei Segmentation

### v2

@@ Abstract
Domain-generalized nuclei segmentation refers to the generalizability of models to unseen domains based on knowledge learned from source domains and is challenged by various image conditions, cell types, and stain strategies. Recently, the Segment Anything Model (SAM) has made great success in universal image segmentation by interactive prompt modes (e.g., point and box). Despite its strengths, the original SAM presents limited adaptation to medical images. Moreover, SAM requires providing manual bounding box prompts for each object to produce satisfactory segmentation masks, so it is laborious in nuclei segmentation scenarios. To address these limitations, we propose a domain-generalizable framework for nuclei image segmentation, abbreviated to NuSegDG. Specifically, we first devise a Heterogeneous Space Adapter (HS-Adapter) to learn multi-dimensional feature representations of different nuclei domains by injecting a small number of trainable parameters into the image encoder of SAM. To alleviate the labor-intensive requirement of manual prompts, we introduce a Gaussian-Kernel Prompt Encoder (GKP-Encoder) to generate density maps driven by a single point, which guides segmentation predictions by mixing position prompts and semantic prompts. Furthermore, we present a Two-Stage Mask Decoder (TSM-Decoder) to effectively convert semantic masks to instance maps without the manual demand for morphological shape refinement. Based on our experimental evaluations, the proposed NuSegDG demonstrates state-of-the-art performance in nuclei instance segmentation, exhibiting superior domain generalization capabilities. The source code is available at https://github.com/xq141839/NuSegDG.
{IEEEkeywords}
Nuclei image segmentation, foundation model fine-tuning, domain generalization
@@ 1 Introduction
Nuclei images are commonly obtained by various imaging modalities, including histopathology slides, fluorescence microscopy, and cryo-electron microscopy. The segmentation task based on such images is crucial for disease diagnosis and treatment planning [1, 2]. In particular, semantic segmentation can be used to calculate the disease area. Instance segmentation aims to identify each nuclear as a separate entity within an image, allowing detailed morphological studies and advanced cellular analysis, such as cell counting. However, the inherent heterogeneity of different modalities, intricate tissue structures and tight cell clustering pose challenges in building a universal nuclei segmentation framework [3, 4, 5, 6].
@@
Traditional U-shape architectures adopt Convolutional Neural Network (CNN) for feature extraction and combine the predicted nuclear proxy maps with morphological post-processing methods to generate instance maps from the semantic segmentation masks [7, 8, 9]. Despite these task-specific models displaying acceptable performance on the seen data, they are difficult to handle unseen domains, especially for the nuclei with different shapes and stain environments. This is because morphological operations are sensitive to the intensity distribution, unexpected artifacts, and noise. This highlights that the generalized nuclei segmentation methods should reduce the dependence on classical image processing algorithms.
@@
The recent emergence of the Segment Anything Model (SAM) [10] has revolutionized segmentation tasks, offering versatile capabilities that surpass traditional methods. SAM has demonstrated exceptional generalization performance in natural image segmentation, showcasing robustness and adaptability across various scenarios [11, 12]. Based on this success, SAM has been applied to a range of medical imaging tasks, revealing its potential to handle diverse and complex segmentation challenges in the medical field, including organ and tissue segmentation and detecting various pathological conditions. These advancements underscore that SAM is promising to provide a robust and generalized solution for diverse medical image segmentation tasks [13, 14, 15, 16]. Despite these advantages, globally fine-tuning SAM requires a large number of pixel-level annotated labels, so it is expensive and impractical for medical scenarios, especially for the specific disease or segmentation task.
@@
Furthermore, SAM mainly adopts interactive prompt modes (e.g., point and box) to guide the segmentation decoding. Although the box mode enables SAM to provide accurate segmentation masks, it is sensitive to the precision of manual annotations and is labor-intensive in nuclei segmentation tasks as each nuclei image usually contains hundreds of cells. On the other hand, the point model is labor-saving, which asks users to click the desired segmentation area. However, current studies have proven that using only one positive point of every cell as the prompt is difficult to drive SAM predicting satisfactory segmentation masks [17, 18]. Therefore, the point prompt mode should be further optimized in nuclei segmentation tasks.
@@
To address these limitations in nuclei image segmentation, we propose a domain-generalizable framework for semantic segmentation and automatic instance map conversion, abbreviated to NuSegDG. It is comprised of three modules: a Heterogeneous Space Adapter (HS-Adapter), a Gaussian-Kernel Prompt Encoder (GKP-Encoder) and a Two-Stage Mask Decoder (TSM-Decoder). Specifically, HS-Adapter is used to adapt the feature representation of SAM from natural to different nuclei images. GKP-Encoder utilizes a single-point prompt to generate the density map with sufficient semantic information for guiding segmentation predictions. TSM-Decoder is responsible for predicting precise semantic segmentation masks and converting them to instance maps without manual morphological image processing. The contributions of this work are summarized as follows:
@@
  • We devise the HS-Adapter that utilizes heterogeneous space projection to adaptively adjust the feature representation of the image encoder based on different nuclei domains, seamlessly harmonizing knowledge transfer between natural and nuclei images.
@@
  • To improve the efficiency of the single-point prompt mode, we introduce the GKP-Encoder that leverages the Gaussian kernel to produce a density map with sufficient semantic prompt information for guiding segmentation predictions.
  • We devise the TSM-Decoder for segmentation decoding. To avoid manual morphological shape refinement, the TSM-Decoder focuses on accurate semantic segmentation mask predictions and directly transforms them into instance maps.
@@
  • We take HS-Adapter, GKP-Encoder, and TSM-Decoder to establish our NuSegDG framework. We conduct extensive experiments on diverse nuclei image datasets, demonstrating that NuSegDG performs better than classical nuclei segmentation methods and state-of-the-art medical SAMs with superior domain generalization capabilities.
@@ 2.1 Nuclei Image Segmentation
The segmentation of nuclei in histopathology images plays an essential role in pathological analysis, enabling pathologists to make precise diagnoses [19]. It can be mainly divided into nuclei semantic segmentation and nuclei instance segmentation. The semantic segmentation focuses on the accuracy of pixel-level classification in each nuclei image, where U-Net [20] has made great success in this task. Over the last decade, researchers mainly focused on improving its ability of feature extraction. Early CNN series leverages the advantages of inductive bias to provide sufficient prior knowledge for accelerating model convergence [21, 22, 23]. Vision Transformer (ViT) [24] further increases the model capacity by utilizing a self-attention mechanism to capture long-range dependencies [25, 26, 27, 28]. The recent Mamba-based frameworks adopted State Space Model (SSM) to optimize the computation complexity of global context [29].
In addition, the instance segmentation task aims to identify each nucleus as a distinct entity. Existing methods usually predict different types of nuclear proxy maps to synthesize instance maps. HoVer-Net [7], Cellpose [30, 31] and CellViT [9], for instance, employed horizontal and vertical distance maps to accurately delineate the boundaries of individual nuclear in histopathology images. CDNet [32] and PROnet [33] respectively utilized directional feature maps and offset maps to enhance the delineation of nuclei boundaries. TSFD-Net [34] and CPP-Net [8] additionally used boundary maps as auxiliary supervisions. Despite their advancements, these methods often require complex post-processing, such as manual morphology operations and thresholding algorithms, to manually synthesize instance maps, so they hinder the generalization capability of models to unseen domains [13]. Our proposed NuSegDG framework addresses these limitations by converting fundamental semantic segmentation masks to instance maps automatically, thereby demonstrating outstanding domain generalization performance across diverse nuclei image domains.
@@ 2.2 SAM for Generalized Medical Image Segmentation
The generalizability of neural networks is crucial for medical image segmentation [35, 36]. Existing methods mainly utilize multi-source domain adaptation [37] and federal learning [38] to address the Domain Generalization (DG) problem. The recent Segment Anything Model (SAM) [10] is a novel interactive architecture that leverages both sparse prompts (e.g., point, box and text) and dense prompts (e.g., mask) to guide the prediction of segmentation masks. Due to its large image encoder, SAM demonstrates robust feature extraction capabilities, enabling outstanding zero-shot generalization across diverse natural image segmentation tasks. On this basis, current studies have explored the potential of SAM in medical image segmentation tasks [39]. For instance, MedSAM [15] and SAMMI [16] globally fine-tuned SAM on more than 10 medical visual modality datasets, achieving notable generalization capabilities with bounding box prompts. However, globally fine-tuning SAM is computationally intensive and needs sufficient training samples due to its large ViT encoder, which is not efficient for nuclei image segmentation tasks. To address this issue, Parameter-Efficient Fine-Tuning (PEFT) techniques have received the most attention from researchers. Methods such as Low-Rank Adaptation (LoRA) [40] and Conv-LoRA [41] injected a set of trainable low-rank matrices into the attention layer of ViT to update the feature representation. Adapter [42] is another common approach used to fine-tune the foundation model [43, 44]. Although these methods reveal their power in homogeneous domain generalization tasks, nuclei images in different domains have disjoint label spaces. Our approach utilizes heterogeneous space mapping to harmonize the feature representation of SAM between natural and nuclei images.
Moreover, various SAM models [15, 17] have demonstrated the necessity of using bounding boxes as prompts to achieve optimal segmentation results in medical imaging. Conversely, relying on single-point prompts often fails to provide sufficient contextual information for accurate segmentation, especially in complex and dense nuclei images [16, 18]. To overcome the limitations of single-point prompts, existing studies introduced extra units, such as YOLO-NAS [45] and GroundingDino [46], to generate prompts. They perform object detection to identify points or bounding boxes, which are then used as prompts for SAM. However, due to the heterogeneity of nuclei images across different datasets, these single-task models often struggle to provide correct prompts, leading to sub-optimal segmentation results. On the contrary, our NuSegDG uses labor-saving single-point annotation to generate sufficient position and semantic prompts, enhancing the generalization capability.
3 Methodology
@@ 3.2 Heterogeneous Space Adapter
Recent studies [12, 16] have highlighted the impressive generalized segmentation capabilities of SAM [10], facilitated by its large-scale image encoder. Especially, the conventional Adapter [42] and LoRA [40] have been widely used to adapt SAM to medical image segmentation [43, 47]. However, such homogeneous space mapping methods are difficult to learn heterogeneous relationships [48] between different nuclei domains. To tackle the issue, we propose the HS-Adapter that leverages heterogeneous space integration to enhance the domain-specific feature representation of nuclei images. Specifically, the input image is first converted into a set of 2D patch embeddings $\mu\subset\mathbb{R}^{(\frac{H\times W}{n})\times d}$, where $H$ and $W$ are height and width of the image, $n=16\times 16$ and $d=768$ stand for the patch size and channels of each patch embedding, respectively. To improve the information interaction within Multi-Head Attention (MHA) layers, the HS-Adapter respectively concatenates learnable parameters $W_{\rm que}=\{(E^{i}_{\rm que},\theta^{i}_{\rm que})\}_{i=1}^{N}$ and $W_{\rm val}=\{(E^{i}_{\rm val},\theta^{i}_{\rm val})\}_{i=1}^{N}$ with the query $\mathcal{Q}$ and value $\mathcal{V}$ branches of SAM, where $E^{i}_{\rm que}$ and $E^{i}_{\rm val}$ are projection layers that map embeddings $\mu$ into feature spaces with $i$-th target mapping channel, $\theta^{i}_{\rm que}$ and $\theta^{i}_{\rm val}$ are up-projections. Additionally, we place the softmax operation $\delta$ on $\mu$ to calculate the weight of each feature space. Finally, $N$ weighted different feature spaces are merged into a heterogeneous space that is used to update the original query and value projection layers of SAM, guiding the computation of attention maps as:
@@ 3.3 Gaussian-Kernel Prompt Encoder
The original SAM [10] and medical SAMs [43, 15, 16] mainly rely on manual box prompts to guide the model in predicting accurate segmentation masks. Despite its advantages, this prompt mode is sensitive to the localization of boxes. Minor labeling errors can significantly reduce the quality of generating segmentation masks. Therefore, the precise manual box annotation is impractical in nuclei segmentation tasks as a histopathological image usually contains thousands of nuclei and tight cell clusters in clinical scenarios. In this paper, we introduce the GKP-Encoder that leverages single-point prompts to produce a high-quality density map, providing additionally sufficient semantic information prompts to assist segmentation decoding.
Given $L$ cell positions: $\{(x_{l},y_{l})\}_{l=1}^{L}$ in a nuclei image, where $x_{l},y_{l}\in\mathbb{N}$, the corresponding density map $\mathcal{D}=\{\mathcal{D}_{z,j}\}\in\mathbb{R}^{H\times W}$ [49] is defined by:

### v3

@@ Abstract
Domain-generalized nuclei segmentation refers to the generalizability of models to unseen domains based on knowledge learned from source domains and is challenged by various image conditions, cell types, and stain strategies. Recently, the Segment Anything Model (SAM) has made great success in universal image segmentation by interactive prompt modes (e.g., point and box). Despite its strengths, the original SAM presents limited adaptation to medical images. Moreover, SAM requires providing manual bounding box prompts for each object to produce satisfactory segmentation masks, so it is laborious in nuclei segmentation scenarios. To address these limitations, we propose a domain-generalizable framework for nuclei image segmentation, abbreviated to NuSegDG. Specifically, we first devise a Heterogeneous Space Adapter (HS-Adapter) to learn multi-dimensional feature representations of different nuclei domains by injecting a small number of trainable parameters into the image encoder of SAM. To alleviate the labor-intensive requirement of manual prompts, we introduce a Gaussian-Kernel Prompt Encoder (GKP-Encoder) to generate density maps driven by a single point, which guides segmentation predictions by mixing position prompts and semantic prompts. Furthermore, we present a Two-Stage Mask Decoder (TSM-Decoder) to effectively convert semantic masks to instance maps without the manual demand for morphological shape refinement. Based on our experimental evaluations, the proposed NuSegDG demonstrates state-of-the-art performance in nuclei semantic and instance segmentation, exhibiting superior domain generalization capabilities. The source code is available at https://github.com/xq141839/NuSegDG.
keywords:
Nuclei segmentation , foundation model , parameter-efficient fine-tuning , domain generalization
journal: Knowledge-Based Systems
@@ 1 Introduction
Nuclei images are commonly obtained by various imaging modalities, including histopathology slides, fluorescence microscopy, and cryo-electron microscopy. The segmentation task based on such images is crucial for disease diagnosis and treatment planning [iqbal2023ldmres, xie2022dmcgnet]. In particular, semantic segmentation can be used to calculate the disease area. Instance segmentation aims to identify each nuclear as a separate entity within an image, allowing detailed morphological studies and advanced cellular analysis, such as cell counting. However, the inherent heterogeneity of different modalities, intricate tissue structures and tight cell clustering pose challenges in building a universal nuclei segmentation framework [caicedo2019nucleus, kumar2017dataset, naylor2018segmentation, mahbod2021cryonuseg].
@@
Traditional U-shape architectures adopt Convolutional Neural Network (CNN) for feature extraction and combine the predicted nuclear proxy maps with morphological post-processing methods to generate instance maps from the semantic segmentation masks [graham2019hover, chen2023cpp, horst2024cellvit]. Despite these task-specific models displaying acceptable performance on the seen data, they are difficult to handle unseen domains, especially for the nuclei with different shapes and stain environments. This is because morphological operations are sensitive to the intensity distribution, unexpected artifacts, and noise. This highlights that the generalized nuclei segmentation methods should reduce the dependence on classical image processing algorithms.
@@
The recent emergence of the Segment Anything Model (SAM) [kirillov2023segment] has revolutionized segmentation tasks, offering versatile capabilities that surpass traditional methods. SAM has demonstrated exceptional generalization performance in natural image segmentation, showcasing robustness and adaptability across various scenarios [zhang2024uv]. Based on this success, SAM has been applied to a range of medical imaging tasks, revealing its potential to handle diverse and complex segmentation challenges in the medical field, including organ and tissue segmentation and detecting various pathological conditions. These advancements underscore that SAM is promising to provide a robust and generalized solution for diverse medical image segmentation tasks [shui2023unleashing, cheng2023sam, ma2024segment, huang2024segment]. Despite these advantages, globally fine-tuning SAM requires a large number of pixel-level annotated labels, so it is expensive and impractical for medical scenarios, especially for the specific disease or segmentation task.
@@
Furthermore, SAM mainly adopts interactive prompt modes (e.g., point and box) to guide the segmentation decoding. Although the box mode enables SAM to provide accurate segmentation masks, it is sensitive to the precision of manual annotations and is labor-intensive in nuclei segmentation tasks as each nuclei image usually contains hundreds of cells. On the other hand, the point model is labor-saving, which asks users to click the desired segmentation area. However, current studies have proven that using only one positive point of every cell as the prompt is difficult to drive SAM predicting satisfactory segmentation masks [zhang2024segment, meng2024nusea]. Therefore, the point prompt mode should be further optimized in nuclei segmentation tasks.
@@
To address these limitations in nuclei image segmentation, we propose a domain-generalizable framework for semantic segmentation and automatic instance map conversion, abbreviated to NuSegDG. It is comprised of three modules: a Heterogeneous Space Adapter (HS-Adapter), a Gaussian-Kernel Prompt Encoder (GKP-Encoder) and a Two-Stage Mask Decoder (TSM-Decoder). Specifically, HS-Adapter is used to adapt SAM from natural to different nuclei images and provides domain-specific feature representations by heterogeneous space integration. GKP-Encoder utilizes a single-point prompt to generate the density map with sufficient semantic information for guiding segmentation predictions. TSM-Decoder is responsible for predicting precise semantic segmentation masks and converting them to instance maps without manual morphological image processing. To the best of our knowledge, we are the first attempting to discover the impact of heterogeneous space integration in domain-generalized nuclei image segmentation during the fine-tuning stage. Secondly, we innovatively leverage Gaussian kernel transforms single-point annotations into high-quality density maps, providing rich semantic and positional prompts. Finally, compared to current state-of-the-art methods [stringer2021cellpose, chen2023cpp, horst2024cellvit], we first adopt a novel semantic-to-instance sequence decoding paradigm to generate final instance segmentation maps, significantly reduce the prediction complexity, and improve generalization capabilities. The contributions of this work are summarized as follows:
  • 1.
@@
    We introduce the HS-Adapter to seamlessly harmonize knowledge transfer between natural and nuclei images and adaptively adjust the feature representation based on different nuclei domains by leveraging heterogeneous space projection.
  • 2.
@@
    We devise the GKP-Encoder that utilizes the labor-saving single-point prompt and Gaussian kernel to produce a density map with sufficient semantic prompt information for guiding segmentation predictions.
  • 3.
    We design the TSM-Decoder that provides a novel semantic-to-instance sequence decoding paradigm to eliminate manual morphological shape refinement and achieve the prediction of accurate instance segmentation maps.
  • 4.
@@
    Our NuSegDG framework integrates HS-Adapter, GKP-Encoder, and TSM-Decoder. We conduct extensive experiments on diverse nuclei image datasets, demonstrating that NuSegDG outperforms classical nuclei segmentation methods and state-of-the-art medical SAM variants and revealing superior domain generalization capabilities.
@@ 2.1 Nuclei Image Segmentation
The segmentation of nuclei in histopathology images plays an essential role in pathological analysis, enabling pathologists to make precise diagnoses [chen2024sam]. It can be mainly divided into nuclei semantic segmentation and nuclei instance segmentation. The semantic segmentation focuses on the accuracy of pixel-level classification in each nuclei image, where U-Net [ronneberger2015u] has made great success in this task. Over the last decade, researchers mainly focused on improving its ability of feature extraction. Early CNN series leverages the advantages of inductive bias to provide sufficient prior knowledge for accelerating model convergence [xu2023dcsau]. Vision Transformer (ViT) [dosovitskiy2021an] further increases the model capacity by utilizing a self-attention mechanism to capture long-range dependencies [lin2022ds, li2023lvit, chen2024transunet]. The recent Mamba-based frameworks adopted State Space Model (SSM) to optimize the computation complexity of global context [ma2024u].
In addition, the instance segmentation task aims to identify each nucleus as a distinct entity. Existing methods usually predict different types of nuclear proxy maps to synthesize instance maps. HoVer-Net [graham2019hover], Cellpose [stringer2021cellpose] and CellViT [horst2024cellvit], for instance, employed horizontal and vertical distance maps to accurately delineate the boundaries of individual nuclear in histopathology images. PROnet [nam2023pronet] utilized offset maps to enhance the delineation of nuclei boundaries. TSFD-Net [ilyas2022tsfd] and CPP-Net [chen2023cpp] additionally used boundary maps as auxiliary supervisions. Despite their advancements, these methods often require complex post-processing, such as manual morphology operations and thresholding algorithms, to manually synthesize instance maps, so they hinder the generalization capability of models to unseen domains [shui2023unleashing]. Our proposed NuSegDG framework addresses these limitations by converting fundamental semantic segmentation masks to instance maps automatically, thereby demonstrating outstanding domain generalization performance across diverse nuclei image domains.
@@ 2.2 SAM for Generalized Medical Image Segmentation
The generalizability of neural networks is crucial for medical image segmentation [li2022domain]. Existing methods mainly utilize multi-source domain adaptation [hu2022domain] and federal learning [liu2021feddg] to address the Domain Generalization (DG) problem. The recent Segment Anything Model (SAM) [kirillov2023segment] is a novel interactive architecture that leverages both sparse prompts (e.g., point, box and text) and dense prompts (e.g., mask) to guide the prediction of segmentation masks. Due to its large image encoder, SAM demonstrates robust feature extraction capabilities, enabling outstanding zero-shot generalization across diverse natural image segmentation tasks. On this basis, current studies have explored the potential of SAM in medical image segmentation tasks. For instance, MedSAM [ma2024segment] and SAMMI [huang2024segment] globally fine-tuned SAM on more than 10 medical visual modality datasets, achieving notable generalization capabilities with bounding box prompts. However, globally fine-tuning SAM is computationally intensive and needs sufficient training samples due to its large ViT encoder, which is not efficient for nuclei image segmentation tasks. To address this issue, Parameter-Efficient Fine-Tuning (PEFT) techniques have received the most attention from researchers. Methods such as Low-Rank Adaptation (LoRA) [hu2021lora] and Conv-LoRA [zhong2024convolution] injected a set of trainable low-rank matrices into the attention layer of ViT to update the feature representation. Adapter [houlsby2019parameter] is another common approach used to fine-tune the foundation model [wu2023medical, lin2023samus]. Although these methods reveal their power in homogeneous domain generalization tasks, nuclei images in different domains have disjoint label spaces. Our approach utilizes heterogeneous space mapping to harmonize the feature representation of SAM between natural and nuclei images.
Moreover, various SAM models [ma2024segment, zhang2024segment] have demonstrated the necessity of using bounding boxes as prompts to achieve optimal segmentation results in medical imaging. Conversely, relying on single-point prompts often fails to provide sufficient contextual information for accurate segmentation, especially in complex and dense nuclei images [huang2024segment, meng2024nusea]. To overcome the limitations of single-point prompts, existing studies introduced extra units, such as YOLO-NAS [terven2023comprehensive] and GroundingDino [liu2023grounding], to generate prompts. They perform object detection to identify points or bounding boxes, which are then used as prompts for SAM. However, due to the heterogeneity of nuclei images across different datasets, these single-task models often struggle to provide correct prompts, leading to sub-optimal segmentation results. On the contrary, our NuSegDG uses labor-saving single-point annotation to generate sufficient position and semantic prompts, enhancing the generalization capability.
3 Method
@@ 3.2 Heterogeneous Space Adapter
Recent studies [zhang2024uv, huang2024segment] have highlighted the impressive generalized segmentation capabilities of SAM [kirillov2023segment], facilitated by its large-scale image encoder. Especially, the conventional Adapter [houlsby2019parameter] and LoRA [hu2021lora] have been widely used to adapt SAM to medical image segmentation [wu2023medical, zhang2023customized]. However, such homogeneous space mapping methods are difficult to learn heterogeneous relationships [huang2024hst] between different nuclei domains. To tackle the issue, we propose the HS-Adapter that leverages heterogeneous space integration to enhance the domain-specific feature representation of nuclei images. Specifically, the input image is first converted into a set of 2D patch embeddings $\mu\subset\mathbb{R}^{(\frac{H\times W}{n})\times d}$, where $H$ and $W$ are height and width of the image, $n=16\times 16$ and $d=768$ stand for the patch size and channels of each patch embedding, respectively. To improve the information interaction within Multi-Head Attention (MHA) layers, the HS-Adapter respectively concatenates learnable parameters $W_{\rm que}=\{(E^{i}_{\rm que},\theta^{i}_{\rm que})\}_{i=1}^{N}$ and $W_{\rm val}=\{(E^{i}_{\rm val},\theta^{i}_{\rm val})\}_{i=1}^{N}$ with the query $\mathcal{Q}$ and value $\mathcal{V}$ branches of SAM, where $E^{i}_{\rm que}$ and $E^{i}_{\rm val}$ are projection layers that map embeddings $\mu$ into feature spaces with $i$-th target mapping channel, $\theta^{i}_{\rm que}$ and $\theta^{i}_{\rm val}$ are up-projections. Additionally, we place the softmax operation $\delta$ on $\mu$ to calculate the weight of each feature space. Finally, $N$ weighted different feature spaces are merged into a heterogeneous space that is used to update the original query and value projection layers of SAM, guiding the computation of attention maps as:
@@ 3.3 Gaussian-Kernel Prompt Encoder
The original SAM [kirillov2023segment] and medical SAMs [wu2023medical, ma2024segment, huang2024segment] mainly rely on manual box prompts to guide the model in predicting accurate segmentation masks. Despite its advantages, this prompt mode is sensitive to the localization of boxes. Minor labeling errors can significantly reduce the quality of generating segmentation masks. Therefore, the precise manual box annotation is impractical in nuclei segmentation tasks as a histopathological image usually contains thousands of nuclei and tight cell clusters in clinical scenarios. In this paper, we introduce the GKP-Encoder that leverages single-point prompts to produce a high-quality density map, providing additionally sufficient semantic information prompts to assist segmentation decoding.
Given $L$ cell positions: $\{(x_{l},y_{l})\}_{l=1}^{L}$ in a nuclei image, where $x_{l},y_{l}\in\mathbb{N}$, the corresponding density map $\mathcal{D}=\{\mathcal{D}_{z,j}\}\in\mathbb{R}^{H\times W}$ [xu2023sppnet] is defined by: